# Weak Adversarial Neural Pushforward Method for Fractional Fokker-Planck Equations

### 13 Mar 2026 v1

@@ 3.3 Monte Carlo Discretization of the Weak Form
where $\bm{\eta}^{(k)}=\{w^{(k)},\kappa^{(k)},b^{(k)}\}$ are the test function parameters.
@@ 3.4 Adversarial Min-Max Training
The generator (pushforward network) is updated by gradient descent to minimize the loss, while the adversary (test function parameters) is updated by gradient ascent to maximize it, thereby searching for the plane-wave modes that are hardest to satisfy. The adversarial update provides an adaptive basis that concentrates on the frequency regime where the current learned distribution violates the weak form most severely.
@@
2:Initialize: pushforward network $F_{\bm{\vartheta}}$; test function parameters $\{w^{(k)},\kappa^{(k)},b^{(k)}\}_{k=1}^{K}$
3:Optimizers: Adam ($\mathrm{lr}=10^{-3}$) for $\bm{\vartheta}$}; Adam ($\mathrm{lr}=10^{-2}$) for $\bm{\eta}$
@@
6:  Compute $\mathcal{L}$ via (18)
7:  ${\bm{\vartheta}}\leftarrow{\bm{\vartheta}}-\nabla_{\bm{\vartheta}}\mathcal{L}$ $\triangleright$ minimize over generator
8:  if epoch $\bmod 10=0$ then
9:   for $j=1$ to 5 do
10:     Recompute $\mathcal{L}$
11:     $\bm{\eta}\leftarrow\bm{\eta}+\nabla_{\bm{\eta}}\mathcal{L}$ $\triangleright$ maximize over test functions
12:   end for
13:  end if
14:end for
@@ 4.1 Symmetric $\alpha$-Stable Lévy Increments
The characteristic function of a symmetric $\alpha$-stable distribution with unit scale is $e^{-|\xi|^{\alpha}}$. Over a time step $\Delta t$, the SDE increment $\Delta L^{(\alpha)}$ has scale $(\Delta t)^{1/\alpha}$. We generate increments using the Chambers-Mallows-Stuck algorithm [4] with $\beta=0$ (symmetry) and scale $(\Delta t)^{1/\alpha}$, implemented via scipy.stats.levy_stable.rvs.
The Euler-Maruyama discretization of (1) reads
@@
with time step $\Delta t=0.01$ and $N_{\mathrm{particles}}=10{,}000$.
Remark 4 (Variance and heavy tails).
For $\alpha<2$, symmetric $\alpha$-stable distributions have infinite variance. In finite-sample particle simulations, the empirical variance fluctuates wildly from run to run due to rare but large excursions. This makes the standard deviation an unreliable comparison metric for $\alpha<2$, as is clearly visible in our results. We therefore rely on robust statistics—median, IQR, MAD, and percentile bands—for all quantitative comparisons.
@@ 5 Numerical Results
5.1 Problem Setup
@@
We solve (2) with the quadratic drift (3) ($k=1$) on $[0,T]$ with $T=2.0$ and $\alpha=1.5$. The initial condition is a Gaussian centered at $x_{0}=1.0$ with standard deviation $\sigma_{0}=0.3$,
$\rho_{0}(x)=\mathcal{N}(x;\,1.0,\;0.09).$ (21)
Figure 1 illustrates the confining potential $V(x)=\tfrac{1}{2}x^{2}$ and the initial density $\rho_{0}$. The initial distribution is centered at $x=1.0$, offset from the potential minimum at $x=0$, so that the drift pushes the distribution toward the origin while the fractional diffusion spreads it with algebraically decaying tails.
@@
Figure 1: Quadratic confining potential $V(x)=\tfrac{1}{2}kx^{2}$ (red, left axis) and initial condition $\rho_{0}=\mathcal{N}(1.0,0.09)$ (blue, right axis) for the fractional Fokker-Planck experiment with $\alpha=1.5$, $k=1.0$. The initial mass is offset from the potential minimum at $x=0$; the dynamics transport it toward the origin while fractional diffusion develops heavy tails.
5.2 Training Configuration
The pushforward network follows the architecture (8) with base dimension $d=5$ and three hidden layers of width 128. The initial distribution samples $x_{0}\sim\mathcal{N}(1.0,0.09)$ are fed as the anchor point, and the latent noise $\mathbf{r}\sim\mathcal{N}(0,I_{5})$ provides stochastic variability.
@@ 5.2 Training Configuration
We use K=2, [^0] K=2{,}000 plane-wave test functions (9), with spatial frequencies $w^{(k)}$ initialized from $\mathcal{N}(0,I)$ and normalized to unit norm, temporal frequencies $\kappa^{(k)}$ initialized from $\mathcal{N}(0,I)$, and phase offsets $b^{(k)}$ from $\mathcal{U}[0,2\pi]$. Batch sizes are M=2, [^0] M=2{,}000 (interior), $M_{0}=M_{T}=1{,}000$ (boundary terms). The adversarial update runs 5 gradient ascent steps every 10 epochs.
Training runs for $N_{\mathrm{epochs}}=1{,}000$ epochs with the Adam optimizer ($\mathrm{lr}=10^{-3}$ for the generator, $\mathrm{lr}=10^{-2}$ for the test functions) and cosine annealing for the generator learning rate. Gradient clipping with maximum norm 1.0 is applied to both sets of parameters.
5.3 Training Convergence
@@
Figure 2 shows the training loss and weak-form residual norm on log-scale axes over 1,000 epochs. Both curves exhibit a sharp initial descent spanning more than two orders of magnitude within the first 50 epochs, followed by a stabilization phase in which the loss oscillates around approximately $10^{-2}$. The oscillatory behavior is characteristic of adversarial min-max training: the test functions periodically find harder frequency modes to probe, causing transient spikes in the loss that the generator then suppresses. The overall convergence is stable and monotonically decreasing on a moving average, confirming that the adversarial dynamics do not lead to mode collapse or divergence.
@@
Figure 2: Training history over 1,000 epochs for the fractional Fokker-Planck problem with $\alpha=1.5$. Left: training loss (18) on a log scale. Right: $\ell^{2}$ norm of the weak-form residual vector $(R^{(1)},\ldots,R^{(K)})$ on a log scale. Both curves exhibit the rapid initial convergence and adversarial oscillation typical of min-max training, with the loss stabilizing around $10^{-2}$ and the residual norm around $10^{1}$ (note that the residual norm is summed over K=2, [^0] K=2{,}000 test functions).
5.4 Transient Distribution Comparison
Figure 3 shows normalized histograms of the neural network samples (blue) alongside the particle simulation histogram (green) at eight evenly-spaced times $t\in\{0.0$, $0.2$, $0.4$, $0.6$, $0.8$, $1.0$, $1.5$, 2.0}2.0\}. At each time, 5,000 samples are drawn from the pushforward network and 10,000 particle simulation trajectories are used.
@@ 5.4 Transient Distribution Comparison
At $t=0$ both histograms accurately reproduce the initial Gaussian centered at $x=1.0$. As time evolves, the distribution drifts toward the potential minimum at $x=0$ and develops the broad, algebraically decaying tails expected from $\alpha$-stable diffusion with $\alpha=1.5$. The neural network distribution tracks this evolution faithfully throughout, capturing both the shift in the bulk of the distribution and the emergence of extended tails. By $t=1.5$ and $t=2$ the distribution has largely relaxed toward the fractional steady state, and the agreement between neural network and particle simulation remains good across the full support visible in the histograms.
@@
Figure 3: Fractional Fokker-Planck evolution ($\alpha=1.5$): neural network samples (blue) versus particle simulation (green) at eight time points. Both histograms are density-normalized over [3,3][-3,3] with 100 bins. The neural network faithfully tracks the drift of the bulk toward the potential minimum at $x=0$ and the development of algebraically heavy tails over time.
5.5 Quantitative Comparison via Robust Statistics
@@
Figure 4 reports six scalar metrics computed from 50 uniformly-spaced time snapshots over $[0,2]$. Four robust metrics—mean, median, IQR, and MAD—show close tracking between neural network and particle simulation throughout the time interval. The median trajectories (upper right panel) are in particularly good agreement, with the neural network median following the particle simulation median through the entire relaxation process. The IQR and MAD curves (center row) both show the progressive widening of the distribution due to fractional diffusion, and the neural network correctly captures the rate of this spreading. The percentile band comparison (lower left) confirms that the 10th-to-90th-percentile spread is well-reproduced, with the neural network generating a distribution whose overall shape and extent match the particle simulation.
@@
The lower right panel displays the standard deviation evolution and illustrates clearly why this metric is inappropriate for $\alpha<2$. The particle simulation standard deviation exhibits large, erratic jumps due to occasional extreme sample values, while the neural network standard deviation is artificially stable because the pushforward map, being bounded during training, does not generate samples in the far tail with sufficient frequency to reproduce the divergent second moment. This contrast underscores the necessity of robust statistics for benchmarking heavy-tailed distributions.

### 20 Mar 2026 v2

@@ 3.4 Monte Carlo Discretisation of the Weak Form
where $\bm{\eta}^{(k)}=\{w^{(k)},\kappa^{(k)},b^{(k)}\}$ are the test function parameters. For the steady-state problem, the same loss is used but with $R^{(k)}=\hat{E}^{(k)}_{\mathrm{ss}}$, the single Monte Carlo estimate of $\mathbb{E}_{\rho_{\infty}}[\mathcal{L}_{\mathrm{ss}}f^{(k)}]$.
@@ 3.5 Adversarial Min-Max Training
The generator (pushforward network) is updated by gradient descent to minimize the loss, improving the pushforward map to satisfy the weak form for the current set of test functions. The adversary (test function parameters) is updated by gradient ascent to maximize the loss, searching for the plane-wave modes that are hardest to satisfy. This adaptive mechanism concentrates the test-function basis on the frequency regime where the current learned distribution most severely violates the weak form.
@@
2:Initialise: pushforward network $F_{\bm{\vartheta}}$; test function parameters $\bm{\eta}=\{w^{(k)},\kappa^{(k)},b^{(k)}\}_{k=1}^{K}$ randomly
3:Optimisers: Adam ($\mathrm{lr}=10^{-3}$) for $\bm{\vartheta}$}; Adam ($\mathrm{lr}=10^{-2}$) for $\bm{\eta}$
@@
6:  Compute $\mathcal{L}({\bm{\vartheta}},\bm{\eta})$ via (24)
7:  Update $F_{\bm{\vartheta}}$: ${\bm{\vartheta}}\leftarrow{\bm{\vartheta}}-\nabla_{\bm{\vartheta}}\mathcal{L}$ $\triangleright$ gradient descent on generator
8:  Update adversarial test functions: $\bm{\eta}\leftarrow\bm{\eta}+\nabla_{\bm{\eta}}\mathcal{L}$ $\triangleright$ gradient ascent on test functions
9:end for
In practice, the learning rate for the generator is decayed using cosine annealing, and gradient clipping with maximum norm 1.0 is applied to both the generator and adversary parameter updates. The two updates may be performed with different frequencies in practice (e.g. multiple adversary steps per generator step) to stabilize the adversarial dynamics, as discussed in [3].
@@ 4.1 Symmetric $\alpha$-Stable Lévy Increments
The characteristic function of a symmetric $\alpha$-stable distribution with unit scale is $e^{-|\xi|^{\alpha}}$. Over a time step $\Delta t$, the SDE increment $\Delta L^{(\alpha)}$ has scale $(\Delta t)^{1/\alpha}$. We generate increments using the Chambers-Mallows-Stuck algorithm [4] with $\beta=0$ (symmetry) and scale $(\Delta t)^{1/\alpha}$, implemented via scipy.stats.levy_stable.rvs. The Euler-Maruyama discretisation of (1) reads
@@
with time step $\Delta t=0.01$ and $N_{\mathrm{particles}}=10{,}000$. For the steady-state benchmarks, we run the SDE for a sufficiently long time $T_{\mathrm{SDE}}=50$ until the empirical distribution has converged to stationarity, as verified by checking that successive time-window histograms agree to within sampling noise.
4.2 Why Robust Statistics Are Required
For $\alpha<2$, symmetric $\alpha$-stable distributions have infinite variance. In finite-sample particle simulations, the empirical variance fluctuates wildly from run to run due to rare but large excursions, making the standard deviation an unreliable comparison metric. Specifically, for a sample of $N$ i.i.d. observations from an $\alpha$-stable law with $\alpha<2$, the sample variance does not converge as $N\to\infty$; instead it grows at rate $N^{2/\alpha-1}$ by the generalized central limit theorem [5].
We therefore rely exclusively on the following robust statistics for all quantitative comparisons in Section 5: the median as a measure of central tendency; the interquartile range (IQR) as a measure of spread; the median absolute deviation (MAD) as a scale estimate; and percentile bands (10th–90th) to summarize the bulk of the distribution. All of these quantities are well-defined and have finite-sample estimators that converge even for $\alpha$-stable distributions with $\alpha<2$.
@@ 5 Numerical Results
We present seven benchmark experiments, all with $\alpha=1.5$. In each case, the pushforward network uses a fully connected architecture with three hidden layers of width 128 and Tanh activations (four layers for the 2D ring experiment). The Adam optimizer is used with learning rate $10^{-3}$ for the generator and $10^{-2}$ for the test functions. The generator learning rate is decayed by cosine annealing, and gradient clipping with maximum norm 1.0 is applied throughout. The experiments are presented in order of increasing difficulty and dimension: we begin with one-dimensional steady-state and transient problems, then extend to two-dimensional steady-state and transient problems, and finally to a five-dimensional transient problem.
Table 1: Summary of numerical experiments ($\alpha=1.5$ throughout). $V$: potential; $T$: time horizon (SDE run time for steady-state benchmarks); $K$: number of test functions; $d$: base dimension; $N_{p}$: particle count. The 2D double-peak potential is given in (29).
Experiment $V$ $T$ $K$ $d$ Epochs $N_{p}$
1D OU (steady) $\tfrac{\theta}{2}(x-\mu)^{2}$ 50 200 5 3000 10000
1D Harmonic (transient) $\tfrac{1}{2}x^{2}$ 2.0 2000 5 1000 10000
1D Double-well (transient) $(x^{2}-1)^{2}$ 2.0 3000 8 10000 10000
1D Triple-well (transient) $x^{2}(x^{2}-1)^{2}$ 2.5 2000 8 5000 10000
2D Double-peak (steady) see (29) 20 (SDE) 300 8 10000 10000
2D Ring (transient) $\tfrac{1}{4}(r^{2}-r_{0}^{2})^{2}$ 0.5 300 8 5000 10000
5D Harmonic (transient) $\tfrac{1}{2}\|\mathbf{x}\|^{2}$ 1.0 2000 5 500 10000
5.1 Experiment 1: Steady-State Ornstein–Uhlenbeck Process
Problem setup.
We consider the fOU process with drift $b(x)=-\theta(x-\mu)$, $\theta=1$, $\mu=2$, and apply the steady-state formulation of Sections 2.22.4. By Proposition 1, the exact steady-state distribution is the symmetric $\alpha$-stable law $S_{1.5}\!\left((2\theta)^{-1/1.5},0,2\right)$ with scale parameter
$c=(2\theta)^{-1/\alpha}=2^{-1/1.5}=2^{-2/3}\approx 0.630.$
This distribution has no closed-form density expression but is characterized entirely by its characteristic function (5). Its tails decay algebraically as $|x-2|^{-2.5}$ by (6), so the distribution has infinite variance, and standard deviation is not an appropriate comparison metric.
Training configuration.
The steady-state pushforward network $G_{\bm{\vartheta}}:\mathbb{R}^{5}\to\mathbb{R}$ has no time input (see Section 3.2). It maps $\mathbf{r}\sim\mathcal{N}(0,I_{5})$ to samples from $\rho_{\infty}$. We use $K=200$ sinusoidal test functions $f^{(k)}(x)=\sin(w^{(k)}x+b^{(k)})$, with spatial frequencies $w^{(k)}$ initialized from $\mathcal{N}(0,1)$ and phase offsets from $\mathcal{U}[0,2\pi]$. The batch size for the Monte Carlo expectation in (10) is M=2, [^0] M=2{,}000 samples. The model is trained for 3,0003{,}000 epochs. The particle benchmark uses 10,000 trajectories of the fOU SDE (26) integrated over $T_{\mathrm{SDE}}=50$ time units at $\Delta t=0.01$.
Results.
Figure 1 shows the learned steady-state distribution alongside the particle benchmark histogram. The learned distribution closely matches the particle histogram across the full displayed support [2,6][-2,6], reproducing both the location of the peak near $\mu=2$ and the asymmetric heavy-tailed character of the particle histogram. The heavier right tail visible in the particle benchmark is consistent with the algebraic decay (6): with 10,000 particles, occasional large excursions skew the empirical histogram, and the learned distribution correctly reproduces the bulk of the distribution while the finite-sample tails fluctuate between runs.
The training loss, shown in the right panel of Figure 1, converges smoothly from an initial value of order $10^{0}$ to below $10^{-2}$ within roughly 500 epochs, after which it stabilizes with the adversarial oscillations characteristic of min-max training. The low number of test functions ($K=200$) and short training (3,0003{,}000 epochs) is sufficient for this one-dimensional steady-state problem, reflecting the simplicity of the objective compared with the transient problems below.
This experiment also serves as a validation of the theoretical result in Proposition 1: by fitting the learned distribution to the known $\alpha$-stable law characterized by (5), one could in principle estimate $\alpha$ and $\theta$ from the pushforward samples without access to the SDE, demonstrating the method’s potential as a tool for statistical inference in fractional diffusion systems.
Figure 1: Steady-state fractional OU ($\alpha=1.5$, $\theta=1$, $\mu=2$): learned distribution (blue, WANPM) versus particle benchmark (orange, 10,000 particles, $T_{\mathrm{SDE}}=50$). The theoretical steady state is the symmetric $\alpha$-stable law $S_{1.5}(2^{-2/3},0,2)$ (Proposition 1) with no closed-form density. The learned distribution faithfully captures the heavy-tailed $\alpha$-stable shape centered at $\mu=2$. Right: training loss on a log scale, 3,000 epochs.
5.2 Experiment 2: Harmonic Confining Potential
Problem setup.
@@ 5.2 Experiment 2: Harmonic Confining Potential
We solve the transient fFPE (2) with the quadratic drift b(x)=kxb(x)=-kx, $k=1$, on the time interval $[0,2]$. The potential is $V(x)=\tfrac{1}{2}x^{2}$, with a single minimum at $x=0$. The initial condition is a Gaussian centered at $x_{0}=1.0$ with standard deviation $\sigma_{0}=0.3$,
$\rho_{0}(x)=\mathcal{N}(x;\,1.0,\;0.09),$ (27)
offset from the potential minimum so that the dynamics are non-trivial: the drift carries the bulk of the distribution toward the origin while the fractional diffusion progressively develops heavy algebraic tails. This is the fractional analog of the classical Ornstein–Uhlenbeck transient problem, for which the $\alpha=2$ solution is a Gaussian with analytically trackable mean and variance. For $\alpha=1.5$, no closed-form solution is available, and we benchmark against particle simulation. Figure 2 shows the potential and the initial density.
@@
Figure 2: Quadratic confining potential $V(x)=\tfrac{1}{2}kx^{2}$ (red, left axis) and initial condition $\rho_{0}=\mathcal{N}(1.0,0.09)$ (blue, right axis), $\alpha=1.5$, $k=1$. The initial mass is offset from the potential minimum at $x=0$; the dynamics transport it toward the origin while fractional diffusion develops algebraically heavy tails.
Training configuration.
@@
The pushforward network uses base dimension $d=5$ and is trained with K=2, [^0] K=2{,}000 plane-wave test functions. The spatial frequencies $w^{(k)}$ are initialized from $\mathcal{N}(0,I)$ and normalized to unit norm; the temporal frequencies $\kappa^{(k)}$ are drawn from $\mathcal{N}(0,I)$; the phase offsets $b^{(k)}$ are sampled from $\mathcal{U}[0,2\pi]$. Batch sizes are M=2, [^0] M=2{,}000 for the interior term and $M_{0}=M_{T}=1{,}000$ for the boundary terms. The model is trained for 1,0001{,}000 epochs with cosine annealing.
Training convergence.
@@
Figure 3 shows the training loss and weak-form residual norm on log-scale axes over the 1,000 training epochs. Both curves exhibit a sharp initial descent spanning more than two orders of magnitude within the first 50 epochs, followed by a stabilisation phase in which the loss oscillates around approximately $10^{-2}$. The oscillatory behavior is characteristic of adversarial min-max training: the test functions periodically find harder frequency modes to probe, causing transient spikes in the loss that the generator then suppresses. The overall convergence is stable and monotonically decreasing on a moving average, confirming that the adversarial dynamics do not lead to mode collapse or divergence.
@@
Figure 3: Training history for the harmonic fFPE ($\alpha=1.5$, 1,000 epochs). Left: training loss (24) on a log scale. Right: $\ell^{2}$ norm of the weak-form residual vector $(R^{(1)},\ldots,R^{(K)})$ on a log scale (summed over K=2, [^0] K=2{,}000 modes). Both stabilize near $10^{-2}$ after rapid initial descent.
Transient distribution comparison.
@@
Figure 4 shows density-normalized histograms of 5,000 pushforward samples (blue) and 10,000 particle-simulation trajectories (green) at eight evenly-spaced times $t\in\{0.0,0.2,0.4,0.6,0.8,1.0,1.5,2.0\}$. Both histograms are normalized over [3,3][-3,3] with 100 bins. At $t=0$ both reproduce the initial Gaussian centered at $x=1.0$ accurately. As time evolves, the distribution drifts toward $x=0$ and develops the characteristic broad, algebraically decaying tails of $\alpha$-stable diffusion. The pushforward network faithfully tracks this evolution throughout, capturing both the shift in the bulk of the distribution and the progressive growth of the tails. By $t=1.5$ and $t=2.0$ the distribution has largely relaxed toward the fractional quasi-steady state, and the agreement between the learned distribution and the particle simulation remains good across the full visible support.
@@
Figure 4: Harmonic potential fFPE ($\alpha=1.5$): pushforward samples (blue) versus particle simulation (green) at eight time points. Both histograms are density-normalized over [3,3][-3,3] with 100 bins.
Quantitative comparison via robust statistics.
@@
Figure 5 reports six scalar metrics computed from 50 uniformly-spaced time snapshots over $[0,2]$. The four robust metrics—mean, median, IQR, and MAD—show close tracking between the pushforward network and the particle simulation throughout the time interval. The median trajectories (upper right panel) are in particularly good agreement, with the learned median closely following the particle simulation median through the entire relaxation process. Both the IQR and MAD curves (center row) correctly capture the progressive widening of the distribution due to fractional diffusion, and the neural network accurately reproduces the rate of this spreading. The percentile band comparison (lower left) confirms that the 10th-to-90th-percentile spread is well reproduced, with the learned distribution matching the particle simulation in overall shape and extent.
@@
The lower right panel of Figure 5 displays the standard deviation evolution and illustrates clearly why this metric is inappropriate for $\alpha<2$. The particle simulation standard deviation exhibits large, erratic jumps due to occasional extreme sample values, while the pushforward map standard deviation is artificially stable because the network does not generate samples in the far tail with sufficient frequency to reproduce the divergent second moment. This contrast underscores the necessity of robust statistics for benchmarking heavy-tailed distributions.